OpenStudy (anonymous):

Given that f(x) = x² + 5x, evaluate f(-2).

OpenStudy (anonymous):

no but do you know 8-4x>-12

OpenStudy (anonymous):

f(-2)=4+5*(-2)=4-10=-6 f(-2)=-6

OpenStudy (tkhunny):

f(0) = 0^2 + 5(0) f(1) = 1^2 + 5(1) f(dog) = dog^2 + 5(dog) f(g(x)) = g(x)^2 + 5(g(x)) Sometimes, you DO just substitute.

OpenStudy (anonymous):

Evaluate g(3) if g(x) = 2x + 1.

OpenStudy (tkhunny):

You do it. You MUST understand this notation. Substitute and simplify.

OpenStudy (anonymous):

i i understood how to do it i would of done it. i wouldnt of posted it on here

OpenStudy (anonymous):

just put the value instead of x

OpenStudy (anonymous):

thanks man

OpenStudy (plainntall):

Given that f(x) = x² + 5x, evaluate f(-2). So what is the value of function at x=-2 (-2)^2+5(-2)=4-10=-6

OpenStudy (tkhunny):

I really do not like that response. Do you see the examples above? What happened? Do you have any course materials? How was it explained? The idea behind this forum is for YOU to show YOUR work so that others may HELP you. Simply doing it for you is not the plan.

OpenStudy (anonymous):

i honestly dont give a flying flutter what its suppose to be for... at all... :D but thanks for waisting your time on typing that long retricestory me me not even to care :D

OpenStudy (tkhunny):

I see. So you get to select the purpose of the forum and not the forum sponsors? Nice. Please read the Code of Conduct. If you wish to live by some other rules, you may wish to do it somewhere else. The slightest amount of social conformity can go a long way. We can't be there on your exams. YOU must learn how to produce the proper results. This takes effort from YOU. Keep in touch.

OpenStudy (anonymous):

idc idc and idc again.... i didnt even read that one